# Evaluating Automatic Speech Recognition Systems for Korean Meteorological Experts

### v2

@@ Abstract
This paper explores integrating Automatic Speech Recognition (ASR) into natural language query systems to improve weather forecasting efficiency for Korean meteorologists. We address challenges in developing ASR systems for the Korean weather domain, specifically specialized vocabulary and Korean linguistic intricacies. To tackle these issues, we constructed an evaluation dataset of spoken queries recorded by native Korean speakers. Using this dataset, we assessed various configurations of a multilingual ASR model family, identifying performance limitations related to domain-specific terminology. We then implemented a simple text-to-speech-based data augmentation method, which improved the recognition of specialized terms while maintaining general-domain performance. Our contributions include creating a domain-specific dataset, comprehensive ASR model evaluations, and an effective augmentation technique. We believe our work provides a foundation for future advancements in ASR for the Korean weather forecasting domain. [^1] Our code and dataset will be publicly available.
@@ 1 Introduction
Meteorological experts utilize diverse data sources for weather forecasting. The database systems they use are vast and complex, making it challenging for those without extensive system knowledge to access the necessary information. Natural language query systems can assist the experts by allowing them to input queries in natural language. The system translates these queries into database queries (e.g., SQL) to search the database (Zhong et al., 2017; Kim et al., 2020; Deng et al., 2022). This process enables forecasters to efficiently access and interpret complex databases without writing complex and structured queries themselves. Notably, Jo et al. (2023) developed an integrated search system for Korean weather data, enabling users to query extensive meteorological information using natural language. This system simplifies data retrieval and thereby improves operational efficiency.
@@
Incorporating Automatic Speech Recognition (ASR) into these systems can improve efficiency by enabling voice-activated queries, which is beneficial for meteorologists who often multitask. However, developing an effective ASR system for the Korean weather domain presents challenges. Most existing ASR models, primarily trained on general domains, may struggle with the specialized vocabulary and pronunciation of meteorological terms (Yadav and Sitaram, 2022). Additionally, despite being trained in multiple languages, many recent multilingual ASR models often underperform in non-English languages, including Korean, due to their unique linguistic characteristics and limited representation in training data (Li et al., 2021; Radford et al., 2023; Ferraz et al., 2024; Song et al., 2024). The agglutinative nature and phonetic complexity of Korean add to these challenges (Lee et al., 2019; Cho et al., 2020). Therefore, the combination of Korean language characteristics and domain-specific vocabulary in meteorology makes developing an effective ASR system particularly difficult.
To investigate the extent of these challenges, we constructed a specialized evaluation dataset to assess how well ASR models perform in the Korean weather domain. We recorded 2,000 spoken queries derived from the natural language dataset created by Jo et al. (2023). Five native Korean speakers recorded the queries, following pronunciation guidelines from domain experts to ensure accurate pronunciation of domain-specific terms. Our evaluation dataset underwent rigorous manual validation to ensure its accuracy and reliability, making it suitable for evaluating ASR systems specifically in the meteorological domain.
Using this dataset, we evaluated popular multilingual ASR models (Radford et al., 2023) in different configurations. Training ASR models with general-domain Korean speech data (Bang et al., 2020) led to decreased performance, highlighting the challenges posed by distributional shifts in specialized domains. For instance, as shown in Fig. 1, both a zero-shot model and a model fine-tuned on a general domain dataset failed to accurately transcribe domain-specific terms (e.g., weather map, average wind speed, and KIM (Korean Integrated Model)) and unit system (e.g., hPa). These results revealed performance issues with domain-specific terminology, underscoring the need for models tailored to this specialized vocabulary. To resolve this issue, we employed a text-to-speech-based data augmentation method, which improved the recognition of specialized terms while maintaining performance on general-domain ASR tasks. We believe our findings provide valuable insights for developing effective domain-specific ASR systems, particularly within the Korean weather forecasting domain. Our contributions can be summarized as follows:
  • We developed a specialized evaluation dataset for Korean weather-related ASR tasks by recording and validating spoken queries from native Korean speakers.
  • We conducted evaluations using different model configurations, identifying areas where performance could be improved with domain-specific dataset creation.
  • We proposed and demonstrated the effectiveness of a text-to-speech-based data augmentation method to improve ASR performance in recognizing specialized meteorological terms.
@@ 2 Related Work
Research on ASR for non-English languages, including Korean, presents unique challenges. Multilingual ASR models, such as Whisper (Radford et al., 2023), often underperform with languages like Korean due to unique linguistic characteristics and limited representation in training datasets (Li et al., 2021; Yadav and Sitaram, 2022; Ferraz et al., 2024; Song et al., 2024). This performance drop is often attributed to the lower proportion and diversity of non-English data in training sets. Efforts to improve ASR for Korean include datasets like KsponSpeech, a spontaneous speech corpus, and ClovaCall, a goal-oriented dialog speech corpus, provide valuable resources for developing and evaluating Korean ASR systems (Bang et al., 2020; Ha et al., 2020). Our research evaluates ASR performance specifically within the Korean meteorological domain, using a tailored evaluation dataset to identify and address domain-specific performance limitations.
@@ 3 Dataset Construction
This study aims to develop a specialized dataset for evaluating Automatic Speech Recognition (ASR) systems within the Korean meteorological domain. The dataset comprises 5,500 spoken queries derived from natural language questions. The construction process includes several key steps: utilizing the natural language question dataset from Jo et al. (2023), obtaining pronunciation guidelines for specialized terms from domain experts, collecting speech data from native speakers, and verifying the correctness of the recorded audio. The detailed data construction pipeline (section 3.1) and the analysis of the resulting dataset (section 3.2) are as follows.
3.1 Data Construction Pipeline
Acquiring Natural Language Questions
We began by utilizing the natural language question dataset released by Jo et al. (2023). This dataset includes a diverse range of questions relevant to the Korean meteorological domain, providing a robust foundation for our spoken query dataset. The dataset consists of two subsets: URL and SQL. We selected 3,575 questions from the URL subset and 1,925 questions from the SQL subset to ensure comprehensive coverage of various meteorological topics that forecasters typically seek. Note that the original dataset in Jo et al. (2023) contains natural language queries mapped to their corresponding structured queries (e.g., SQL). For our purposes, we used only natural language queries, focusing on how these questions would be articulated by users in real-world scenarios. For more details about natural language questions, please refer to the original paper of the dataset (Jo et al., 2023).
Obtaining Pronunciation Guidelines from Domain Experts
To ensure the accurate pronunciation of specialized meteorological terminology, we collaborated with domain experts. These experts provided comprehensive guidelines on how to pronounce specific terms correctly. For example, the acronym "BUFR" should be pronounced as "buffer" and "AMI" as "A-M-I." Additionally, terms such as "33009 station" could be pronounced as "three three zero zero nine station" or "thirty-three thousand and nine station." These detailed instructions were crucial for maintaining the linguistic integrity of the dataset, ensuring that all domain-specific terms were articulated accurately by the speakers. More detailed guidelines are provided in Appendix A.
@@ 3.1 Data Construction Pipeline
Uttr. time (sec) 7.052.55
- min time 0.92
- max time 29.98
Avg. words 7.594.34
Avg. chars 24.4915.62
@@
Table 1: Dataset Statistics. Uttr. time (sec) denotes the audio lengths. Words are tokenized using the NLTK library (Bird and Loper, 2004). Absent ratio (%) refers to the percentage of unique words in our dataset that are not present in the Korean ASR dataset for open-domain dialog (Bang et al., 2020), divided by the total number of unique words in our dataset. Numbers in subscript represent standard deviation values.
Collecting Speech Data from Native Speakers
Eleven native Korean speakers were recruited to record the spoken queries. Seven were male, and four were female, all in their twenties. Each participant was given the pronunciation guidelines and detailed instructions on how to record the questions. Participants were compensated at a rate of 16,000 KRW (approximately 11.6 USD) per 100 queries. The recordings were conducted in controlled environments to minimize background noise and ensure high audio quality. This phase resulted in a substantial collection of audio recordings that correspond to the natural language questions.
Verifying the Recorded Voices
Following the collection of speech data, a rigorous verification process was implemented. Human verifiers listened to each recording to ensure accuracy and adherence to the pronunciation guidelines. They also checked for clarity and the absence of background noise or errors. Two of the authors are employed for this verification process. This process was essential to guarantee the quality and reliability of the dataset for subsequent ASR evaluation.
@@ 3.2 Dataset Analysis
Table 1 presents the overall statistics of our dataset, which consists of 5,500 spoken queries. The average utterance time is 7.05 seconds, ranging from 0.92 to 29.98 seconds, with each utterance containing an average of 7.59 words and 24.49 characters. The proportion of unique words in our dataset that are not found in the general domain ASR dataset (Bang et al., 2020) is 24.86%, highlighting the specialized and domain-specific nature of our dataset.
Table 2 shows the top 9 most frequently occurring words in both general and weather domain datasets. The general domain contains common conversational words, whereas the weather domain includes specialized terminology, underscoring the necessity of domain-specific datasets for improving ASR performance in meteorological applications.
Table 2: Nine most frequently occurred words in Korean ASR datasets in open-domain dialog (General) and weather (Weather) domains. The KsponSpeech (Bang et al., 2020) is used as a general domain dataset. Each English word is manually translated by the authors from the original Korean word presented in the bracket.
@@ 4.1 Evaluation Datasets
We use our Korean ASR dataset for the weather domain to evaluate different ASR models. The dataset comprises 5,500 spoken queries, which we randomly split into 500 samples for the development set and 5,000 samples for the test set. Additionally, we utilize the eval-clean test set from the KsponSpeech dataset (Bang et al., 2020) to evaluate the models in a general conversation domain. This dataset consists of 3,000 audio and their corresponding answer transcriptions. This approach allows us to compare models across both domain-specific and general ASR tasks.
@@ 4.3 Models and Training Datasets
Zero-shot Multilingual Models
For our experiments, we used the multilingual Whisper model family (Radford et al., 2023), evaluating four different model sizes: tiny (39M), small (244M), medium (769M), and large-v2 (1550M). These pre-trained models were assessed in a zero-shot manner to evaluate their performance without additional fine-tuning on our dataset. The target language and task in Whisper’s prefix tokens were set to Korean and transcribe, respectively. Additionally, for evaluation in the weather domain, we tested Google’s commercial speech-to-text API [^2] https://cloud.google.com/speech-to-text.
Fine-tuning on General Open-domain Dialogues
Besides zero-shot evaluation setup, to explore the impact of adapting a multilingual ASR model to the Korean language, we fine-tuned an ASR model using the KsponSpeech (Bang et al., 2020) dataset. The dataset contains 619k training instances about open-domain dialogue utterances from Native Korean annotators. The Whisper large-v2 is fine-tuned on this dataset.
Data Augmentation for Meteorological Domain
Inspired by Zheng et al. (2021), we hypothesize that teaching the model the pronunciation of specialized terms in the weather domain is crucial for accurate transcription. To facilitate this, we used a simple text-to-speech (TTS) system to generate audio for natural language queries in the weather domain, which was then used for training. We selected 10k and 9.8k instances from the URL and SQL subset of Jo et al. (2023) as the training queries. To avoid test set leakage, the natural language queries in our human-recorded evaluation dataset were excluded from these training queries. Using the Google TTS service [^3] https://gtts.readthedocs.io/, we generated audio for the selected queries. English words in the queries were converted to their Korean pronunciations before being provided to the TTS service based on expert guidelines described in Section 3.1. The resulting TTS-generated audio dataset was used as a training dataset. We fine-tuned the Whisper large-v2 either solely on the TTS-based datasets or on both the TTS-based datasets and the KsponSpeech dataset, resulting in two different ASR models.
Post-processing with Unimodal LLMs
To evaluate the effectiveness of large language models (LLMs) in refining ASR transcriptions without utilizing acoustic information (Chen et al., 2023; Hu et al., 2024a), we conducted an experiment where GPT-4o-mini (Achiam et al., 2023) post-processed the top-1 transcription from Whisper-large-v2. Specifically, we provided the LLM with the single best output from the ASR model. The model was tested in both a zero-shot setting and a 20-shot setup using examples from the development set. We adopted the text prompt from (Chen et al., 2023) and tailored it to the Korean meteorological domain to ensure corrections aligned with domain-specific terminology. (i.e., "Below is the best-hypotheses transcribed from speech recognition system for Korean Meteological experts. Please try to revise it and write the response for the true transcription.") This experiment allows us to assess how effectively an LLM can improve ASR accuracy in a specialized domain when limited to a single-hypothesis input.
@@ 4.4 Implementation Details
All fine-tuned models are parameter-efficiently trained using LoRA (Hu et al., 2022) with $r$=32. The fine-tuning process involved 3 epochs of training with a batch size of 48, a learning rate of 1e-3, and a warmup ratio of 0.1 with AdamW optimizer (Loshchilov and Hutter, 2017). For fine-tuned models, we saved the checkpoint after each epoch and selected the best one based on the lowest CER score from the development set of our dataset. The greedy decoding (Holtzman et al., 2020) is used as a decoding algorithm for all models. All models are implemented with Transformers framework (Wolf et al., 2020) and PyTorch (Paszke et al., 2019).
Figure 3: Effect of LLM-based post-processing on ASR. We compare ASR performance under three conditions: without LLM-based grammatical error correction (w/o GEC), with zero-shot post-processing (0 Shot), and with a 20-shot setup (20 Shot).
@@ 5 Results and Analyses
We present the key findings from our ASR evaluation on the Korean meteorological dataset. Table 3 summarizes the results across different setups.
@@
Zero-shot ASR models usually struggle with domain-specific terminology
@@
Whisper models exhibit high CER and WER when transcribing meteorological queries, particularly for specialized terms. While larger models (e.g., Whisper-large-v2) reduce errors compared to smaller variants, they still struggle with domain-specific terminology, reinforcing the need for adaptation.
@@
Fine-tuning improves general performance, but domain adaptation is crucial
@@
Fine-tuning Whisper-large-v2 on KsponSpeech enhances general transcription accuracy, lowering CER and WER across datasets. However, performance gains on meteorological queries remain limited, suggesting that general-domain adaptation alone is insufficient for specialized terminology.
Synthetic data (TTS) effectively enhances domain adaptation
@@
Training with TTS-generated weather-domain data significantly improves transcription accuracy for meteorological terms. The model fine-tuned on both KspSpeech and synthetic weather data achieves the lowest CER (8.68%) and WER (22.58%) on the test set, demonstrating the effectiveness of controlled synthetic augmentation.
LLM-based post-processing does not consistently improve ASR output
@@
As shown in Figure 3, the LLM-based post-processing does not consistently improve meteorological term recognition. The CER, WER, and sWER remain higher than those of the zero-shot Whisper model, indicating that text-based refinement alone is insufficient for domain-specific transcription without additional acoustic training. While training a specialized GEC model on the target dataset or leveraging large-N hypothesis generation could potentially enhance performance (Chen et al., 2023; Hu et al., 2024b), these approaches would require a large-scale training corpus and significantly increase computational costs during inference.
Commercial ASR systems underperform in specialized tasks
Google’s STT model, a widely used commercial ASR system, yields the highest error rates (e.g., 20.38% on Weather-test set), performing worse than zero-shot Whisper-large-v2. This highlights the limitations of generic ASR systems in handling meteorological speech and underscores the necessity of domain-specific adaptation.
These results emphasize the importance of fine-tuning ASR models with domain-specific data. While general LLMs can aid transcription refinement, substantial performance gains require dedicated acoustic adaptation.
Answer 8월 30일 증기압과 기온 평년값은 뭐니 ZS 8월 30일 증기압과 기운 평량값은 뭐니 FTGeneral 8월 30일 기압과 기온 평년값은 뭐니 FTBoth 8월 30일 증기압과 기온 평년값은 뭐니 Google STT 8월 30일 증기압과 기온 평년 값은 뭐니 LLM20shot 8월 30일 증기압과 기온 평값은 무엇인가요 Answer 관측횟수 CALM 0 5m/s 순별 하순 211 인제 ZS 관측해수컴 0 5mf 순별 하순 211 인 FTGeneral 관측 해수 칸 영 5m 퍼셉커 순별 하순 211 인제11 인제 FTBoth 관측해서 CALM 0 5m순별 하순 2 1 1 인제 Google STT 관측에서 컴 0 5 m 버서커 순별 순 2 인제 LLM20shot 관측 0$\sim$5 m/s 순별 하순 211 인
@@
Case Study
To illustrate common ASR challenges, we examine qualitative examples from Table 4. One issue arises with numerical and meteorological term recognition. In a wind forecast transcription, "지상 시계열 바람장미 풍속계급별 판촉횟수 CALM 0 5m/s 순별 하순 211 인제" ("Ground-level wind rose classified by wind speed, CALM 0–5m/s, late-stage 211 Inje"), zero-shot Whisper and Google STT misinterpret "CALM", a standard meteorological term for low wind speeds, distorting critical numerical values. Another case involves unintended style shifts. In "8월 30일 증가압과 기온 평년값은 뭐니" ("What are the historical averages for pressure and temperature on August 30?"), the post-processed output rephrases it as "8월 30일 증가압과 기온 평년값은 무엇인가요?", making the tone more formal. While this improves fluency, it may be undesirable for contexts requiring colloquial speech. These cases highlight two key challenges: (1) accurately preserving domain-specific terms and numerical values, and (2) balancing grammatical refinement with maintaining the intended speech style. Addressing these requires targeted domain adaptation and context-aware correction strategies.

### v3

@@ Abstract
Automatic speech recognition systems often fail on specialized vocabulary in tasks such as weather forecasting. To address this, we introduce an evaluation dataset of Korean weather queries. The dataset was recorded by diverse native speakers following pronunciation guidelines from domain experts and underwent rigorous verification. Benchmarking both open-source models and a commercial API reveals high error rates on meteorological terms. We also explore a lightweight text-to-speech-based data augmentation strategy, yielding substantial error reduction for domain-specific vocabulary and notable improvement in overall recognition accuracy. Our dataset is available at https://huggingface.co/datasets/ddehun/korean-weather-asr.
@@ 1 Introduction
Meteorologists rely on vast, complex databases for forecasting, yet crafting precise SQL queries demands specialized expertise. Natural language interfaces address this gap by translating user questions into database queries (Zhong et al., 2017; Kim et al., 2020; Deng et al., 2022). Notably, Jo et al. (2023) developed an integrated search system for Korean weather data, enabling users to query extensive meteorological information using natural language. This system simplifies data retrieval and thereby improves operational efficiency.
Incorporating Automatic Speech Recognition (ASR) into these systems can further enhance usability by enabling voice-driven queries, a practical advantage for busy forecasters. However, off-the-shelf ASR models—often trained on general-domain or English-centric corpora—tend to misrecognize Korean meteorology terms due to both the language’s agglutinative structure and specialized vocabulary (Lee et al., 2019; Cho et al., 2020; Li et al., 2021; Yadav and Sitaram, 2022; Radford et al., 2023; Ferraz et al., 2024; Song et al., 2024).
@@
To quantify these challenges, we created an evaluation set of 5,500 Korean weather queries. Eleven native speakers recorded queries sourced from Jo et al. (2023), following expert pronunciation guidelines and a thorough manual validation process to guarantee correctness. We then benchmarked several popular multilingual ASR models (e.g., Whisper variants) and observed consistent misrecognitions of domain-specific terms—such as weather map, average wind speed, KIM (Korean Integrated Model), and units like hPa—even after fine-tuning on a large general-domain Korean speech corpus (Bang et al., 2020) (Fig. 1). We also explored lightweight approaches on our dataset, including TTS-based data augmentation (Zheng et al., 2021) and LLM-based post-processing (Hu et al., 2024a).
Our contributions are as follows: (1) We construct a domain-specific ASR evaluation dataset by recording and validating weather-related queries from native Korean speakers; (2) we evaluate various ASR configurations and highlight the need for domain-specific adaptation; and (3) we explore TTS-based data augmentation to improve recognition of specialized meteorological terms.
@@ 2 Related Work
Research on ASR for non-English languages, including Korean, presents unique challenges. Multilingual ASR models, such as Whisper (Radford et al., 2023), often underperform with languages like Korean due to unique linguistic characteristics and limited representation in training datasets (Li et al., 2021; Yadav and Sitaram, 2022; Ferraz et al., 2024; Song et al., 2024). This performance drop is often attributed to the lower proportion and diversity of non-English data in training sets. Efforts to improve ASR for Korean include datasets like KsponSpeech, a spontaneous speech corpus, and ClovaCall, a goal-oriented dialog speech corpus, provide valuable resources for developing and evaluating Korean ASR systems (Bang et al., 2020; Ha et al., 2020). We evaluate ASR performance specifically within the Korean meteorological domain, using a tailored evaluation dataset to identify and address domain-specific performance limitations.
@@ 3 Dataset Construction
This study aims to develop a specialized dataset for ASR systems within the Korean meteorological domain. The construction process includes several key steps: Acquiring the natural language questions, obtaining pronunciation guidelines for specialized terms from domain experts, collecting speech data from native speakers, and verifying the correctness of the recorded audio. The detailed data construction pipeline (section˜3.1, Fig. 2) and the analysis of the resulting dataset (section˜3.2) are as follows.
@@
Utterance time (sec) 7.052.55
- min / max time 0.92 / 29.98
Avg. chars / words 24.4915.62 / 7.594.34
@@
Table 1: Dataset Statistics. Absent ratio (%) refers to the percentage of unique words not presented in the general Korean ASR dataset (Bang et al., 2020), divided by the total number of unique words in our dataset.
3.1 Data Construction Pipeline
Acquiring Natural Language Questions We used the natural language question dataset released by Jo et al. (2023), which covers diverse meteorological queries. From its URL and SQL subsets, we selected 3,575 and 1,925 questions, respectively, to ensure broad topic coverage. While the original dataset includes structured query mappings (e.g., SQL), we used only the natural language questions, focusing on how forecasters would verbally express their information needs. For more on the original dataset, see Jo et al. (2023).
Obtaining Pronunciation Guidelines from Domain Experts To ensure the correct pronunciation of domain-specific terms, we collaborated with meteorological experts. They provided detailed guidelines—for example, pronouncing "BUFR" as "buffer" and "AMI" as "A-M-I". Additionally, terms such as "33009 station" could be pronounced as "three three zero zero nine station" or "thirty-three thousand and nine station." These guidelines helped speakers produce consistent and intelligible renditions of specialized terms. Full details are in Appendix A.
Table 2: Nine most frequently occurring words in Korean ASR datasets in open-domain dialog (General) and weather (Weather) domains. The KsponSpeech (Bang et al., 2020) is used as a general domain dataset.
Collecting Speech Data from Native Speakers We recruited eleven native Korean speakers (7 male, 4 female, all in their twenties) to record the queries. Participants were provided with the pronunciation guidelines and detailed instructions. Each was compensated at a rate of 16,000 KRW (approx. 11.6 USD) per 100 queries. All recordings were conducted in quiet and controlled environments.
Verifying the Recorded Voices Following the collection of speech data, a rigorous verification process was implemented. Human verifiers listened to each recording to ensure accuracy and adherence to the pronunciation guidelines. They also checked for clarity and the absence of background noise or errors. Two of the authors are employed for this verification process. This process was essential to guarantee the quality and reliability of the dataset for subsequent ASR evaluation.
@@ 3.2 Dataset Analysis
Table 1 shows the overall statistics of our dataset, which includes 5,500 spoken queries. The average utterance length is 7.05 seconds, with an average of 7.59 words and 24.49 characters per query. Notably, 24.86% of the unique words in our dataset are absent from the general-domain ASR corpus (Bang et al., 2020), highlighting the specialized vocabulary of meteorological speech.
Table 2 lists the most frequent words in general-domain and weather-domain datasets. While general speech contains common conversational terms, our dataset prominently features specialized meteorological vocabulary. This contrast underscores the need for domain-specific datasets to improve ASR performance in technical applications.
@@ 4.1 Evaluation Datasets
We use our Korean ASR dataset for the weather domain to evaluate different ASR models. The dataset comprises 5,500 spoken queries, which we randomly split into 500 samples for the development set and 5,000 samples for the test set. Additionally, we utilize the eval-clean test set from the KsponSpeech dataset (Bang et al., 2020) to evaluate the models in a general conversation domain. This dataset consists of 3,000 audio files and their corresponding answer transcriptions.
@@ 4.3 Models and Training Datasets
Zero-shot Multilingual Models For our experiments, we used the multilingual Whisper model family (Radford et al., 2023), evaluating four different model sizes: tiny (39M), small (244M), medium (769M), and large-v2 (1550M). These pre-trained models were assessed in a zero-shot manner to evaluate their performance without additional fine-tuning on our dataset. The target language and task in Whisper’s prefix tokens were set to Korean and transcribe, respectively.
Fine-tuning on General Open-domain Dialogues To explore the impact of adapting a multilingual ASR model to the Korean language, we fine-tuned an ASR model using the KsponSpeech (Bang et al., 2020) dataset. The dataset contains 619k training instances about open-domain dialogue utterances from Native Korean annotators. The Whisper large-v2 is fine-tuned on this dataset. More implementation details are in Appendix 4.4.
Data Augmentation for Meteorological Domain Inspired by Zheng et al. (2021), we hypothesize that teaching the model the pronunciation of specialized weather terms is crucial for accurate transcription. To this end, we used a TTS system to generate audio for 10k and 9.8k natural language queries from the URL and SQL subsets of Jo et al. (2023), respectively. Queries in our evaluation set were excluded to prevent test leakage. We employed the Google TTS service [^1] https://gtts.readthedocs.io/ and converted English words to Korean pronunciations based on expert guidelines (Section 3.1) before synthesis. The resulting TTS audio was used to fine-tune Whisper large-v2, either alone or combined with KsponSpeech, yielding two ASR variants.
Post-processing with Unimodal LLMs To assess whether large language models (LLMs) can improve ASR outputs without acoustic input (Chen et al., 2023; Hu et al., 2024a), we used GPT-4o-mini (Achiam et al., 2023) to revise the top-1 transcription from Whisper-large-v2. The model was evaluated in both zero-shot and 20-shot settings, using examples from the development set. We adapted the prompt from Chen et al. (2023) to the Korean meteorological domain to better handle domain-specific terms. This setup allows us to test how effectively an unimodal LLM can refine ASR transcriptions based solely on textual input.
@@ 4.4 Implementation Details
All fine-tuned models are parameter-efficiently trained using LoRA (Hu et al., 2022) with rr=32. The fine-tuning process involved 3 epochs of training with a batch size of 48, a learning rate of 1e-3, and a warmup ratio of 0.1 with AdamW optimizer (Loshchilov and Hutter, 2017). For fine-tuned models, we saved the checkpoint after each epoch and selected the best one based on the lowest CER score from the development set of our dataset. The greedy decoding (Holtzman et al., 2020) is used as a decoding algorithm for all models. All models are implemented with the Transformers framework (Wolf et al., 2020) and PyTorch (Paszke et al., 2019). For the post-processing with unimodal LLMs, we used the following text prompt: "Below is the best-hypotheses transcribed from speech recognition system for Korean Meteorological experts. Please try to revise it and write the response for the true transcription."
@@ 5 Results and Analyses
In this section, we describe the key observations. Table 3 summarizes results across different setups.
@@
Zero-shot ASR models struggle with domain-specific terminology
@@
Whisper models show high CER and WER when transcribing meteorological queries, particularly for specialized terms. Larger models (e.g., Whisper-large-v2) reduce errors relative to smaller variants but still struggle with domain-specific vocabulary, highlighting the need for further adaptation.
@@
Fine-tuning improves general performance, but domain-specific adaptation is essential
@@
Fine-tuning Whisper-large-v2 on KsponSpeech lowers CER and WER overall, but improvements on meteorological queries are limited. This suggests that general-domain data alone is insufficient for handling specialized terms.
@@
Synthetic data (TTS) aids domain adaptation Training with TTS-generated weather-domain queries improves transcription of meteorological terms. The model fine-tuned on both KsponSpeech and synthetic data achieves balanced performance across general (CER 10.54%) and weather domains (CER 17.46%), while the model trained only on synthetic data achieves the lowest error rates in the weather domain (CER 10.64%).
LLM-based post-processing yields inconsistent gains
@@
As shown in Figure 3, LLM-based post-processing does not consistently improve recognition of domain-specific terms. Error rates remain higher than the zero-shot baseline, suggesting that text-only refinement is insufficient without acoustic adaptation. While training a dedicated GEC model or using N-best hypotheses (Chen et al., 2023; Hu et al., 2024b) may help, these approaches require large corpora and increase inference costs.
Commercial ASR systems perform poorly on specialized data
Google’s STT model yields the highest error rates (e.g., 20.38% on the Weather-test set), underperforming even the zero-shot Whisper-large-v2. This underscores the limitations of generic ASR systems in specialized domains and the need for targeted adaptation. Overall, these results highlight the importance of fine-tuning with domain-specific data. While LLMs can support transcription refinement, substantial improvements require acoustic-level adaptation.
Figure 3: Effect of LLM-based post-processing. We compare three conditions: without LLM-based grammatical error correction (w/o GEC), with zero-shot (0 Shot), and with 20-shot (20 Shot) setups.
Case Study
To illustrate common ASR errors, we present examples from Table 4. One issue involves domain-specific term misrecognition. In a wind forecast transcription, "지상 시계열 바람장미 풍속계급별 판촉횟수 CALM 0 5m/s 순별 하순 211 인제" ("Ground-level wind rose classified by wind speed, CALM 0–5m/s, late-stage 211 Inje"), both zero-shot Whisper and Google STT misinterpret "CALM", distorting critical numerical information. Another case involves unintended style shifts. In "8월 30일 증가압과 기온 평년값은 뭐니" ("What are the historical averages for pressure and temperature on August 30?"), the post-processed output rephrases it more formally as "8월 30일 증가압과 기온 평년값은 무엇인가요?". While fluency improves, such shifts may be undesirable in contexts requiring a colloquial tone. These cases reflect two key challenges: (1) preserving domain-specific and numerical terms, and (2) refining grammar while maintaining intended speech style. Addressing these requires domain-aware and context-sensitive correction strategies.
[…]